Plates OABC with weight 500 N is being held by two ropes that connects to point D. The tension in rope BD and CD each are 310 N and 270 N. Determine an equivalent system of force and coupled about point O and also the angle between the ropes.

I cm À Z | 0,25 cm О, 2 cm
Answer #1


Coordinates in 0 = (0,0,0) C=(0,0.1,0) B = (0,2,0-1,0) D=CO-05, 0, 0.25) G = 0.02, 01,0) = (0.1,0.05,0) Pep To-Te + C005-o+60

The Top U CD= Top (P) CD = 210 0.057-0.13+0.25 R Voos 8+2015+0.25% Tas = 49.2957 - 981595 + 246.476 W = -WR = -500 R Equivale

couple about point o (mo) = Falot Test Palot Foot hoe ra, the ño - S0-17x [49.2957–98.5454 246.476K)+ (0.27 + 0.13 ) x (-150.